OpenStudy (anonymous):

please? Let f(x) = |x - 5|. Which of the following conditional statements hold true? Check all that apply. A. f(x) = -(x - 5) if x < 5 B. f(x) < 0 if x = 0 C. f(x) = x - 5 if x > 0 D. f(x) = x - 5 if x > 5

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@kevsturge

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

f(x) is the distance of x - 5 from 0.

Parth (parthkohli):

Either -(x - 5) or x - 5

OpenStudy (anonymous):

SO A C D ?!? @ParthKohli

OpenStudy (anonymous):

first one is definitely true

OpenStudy (anonymous):

what about c and d?

OpenStudy (anonymous):

omg ur a guru! GIVE ME UR BRAIN ?

Parth (parthkohli):

Lol satellite is the best member on the site.

OpenStudy (anonymous):

c is not true. lets try it with 1 because we agree that \(1>0\) right? \(|1-5|=|-4|=4\) but \(1-5=-4\) so that is false

OpenStudy (anonymous):

@ParthKohli thats awseome! ur awesome too :) <3

Parth (parthkohli):

truedat

OpenStudy (anonymous):

parth is a fast answerer for sure. more accurate than i am

OpenStudy (anonymous):

yea hes so fast! OR SHE?

OpenStudy (anonymous):

wait so do u guys think its a and d? BECAUSE IT CANT BE B RIGHT?

Parth (parthkohli):

Oh thank you satellite :) you are an old member on here so I respect you. I'm just 12 years old so no comparison.

Parth (parthkohli):

Try it with D, @sktjell Absolute value means distance from zero on a number line in a nutshell

OpenStudy (anonymous):

in any case you answers are A and D because \[f(x) = |x-5| = \left\{\begin{array}{rcc} -(x-5) & \text{if} & x < 5 \\ x-5& \text{if} & x \geq 5 \end{array} \right. \]

Parth (parthkohli):

It doesn't state equal to but that still proves it.

OpenStudy (anonymous):

tahnks @satellite73 and omg i cant believe ur 12 @ParthKohli wow i feel even more dumb lol

Parth (parthkohli):

Hmm you shouldn't . Math is God's language. You should study it deeply.

OpenStudy (mathteacher1729):

@sktjell Have you tried using geogebra yet? I'll stop asking now. But my goodness, take it fro me, it is very helpful for the kinds of problems you're asking. :)